# MapGPT: Map-Guided Prompting with Adaptive Path Planning for Vision-and-Language Navigation

### v2

@@ Abstract
Embodied agents equipped with GPT as their brain have exhibited extraordinary decision-making and generalization abilities across various tasks. However, existing zero-shot agents for vision-and-language navigation (VLN) only prompt the GPT-4 to select potential locations within localized environments, without constructing an effective “global-view” for the agent to understand the overall environment. In this work, we present a novel map-guided GPT-based agent, dubbed MapGPT, which introduces an online linguistic-formed map to encourage the global exploration. Specifically, we build an online map and incorporate it into the prompts that include node information and topological relationships, to help GPT understand the spatial environment. Benefiting from this design, we further propose an adaptive planning mechanism to assist the agent in performing multi-step path planning based on a map, systematically exploring multiple candidate nodes or sub-goals step by step. Extensive experiments demonstrate that our MapGPT is applicable to both GPT-4 and GPT-4V, achieving state-of-the-art zero-shot performance on the R2R and REVERIE simultaneously ($\sim$10and $\sim$12improvements in SR), and showcasing the newly emerged global thinking and path planning abilities of the GPT.
@@
[…]
@@ 1 Introduction
Large language models (LLMs) Touvron et al. (2023a, b); Chowdhery et al. (2022); Anil et al. (2023) have demonstrated strong performance in various domains. As the most powerful LLMs, the GPT series models Brown et al. (2020); OpenAI (2023a, b, c) can even serve as the brain of embodied agents Huang et al. (2023); Mu et al. (2023); Ahn et al. (2022), enabling them to engage in explicit thinking and decision-making process. Moreover, these GPT-based agents are typically zero-shot or few-shot, eliminating the burdensome tasks of data annotation and model training, and they also demonstrate remarkable sim-to-real abilities.
@@
However, such zero-shot VLN agents face some challenges. Firstly, these methods are developed for a multi-expert pure language system based on GPT-4, necessitating the conversion of visual observations into text and multiple rounds of text summarizations, which inevitably leads to information loss and hinders the application to multimodal LLMs (e.g., GPT-4V). More importantly, these agents make decisions based solely on the observations of the local environment. As shown in Figure 1 (left), given only the local action space, when an agent realizes that it has engaged in an erroneous exploration, it can only continue to explore surrounding environment aimlessly.
@@ 2 Related Work
Vision-and-Language Navigation (VLN) As a representative multi-modal embodied AI task, VLN requires an agent to combine human instructions and visual observations to navigate and locate targets in real-world scenes. Previous learning-based approaches Wang et al. (2019); Ma et al. (2019); Deng et al. (2020); Qi et al. (2020a) proposed various model architectures and trained their models on domain-specific datasets. Besides, pretrained models Hong et al. (2021); Chen et al. (2021b, 2022); Qiao et al. (2022); Guhur et al. (2021); An et al. (2022); Qiao et al. (2023); Wang et al. (2023); Pan et al. (2023) have been widely applied to produce better multi-modal representations. Recently, to address the reliance on domain-specific data and the possible sim-to-real gap, some zero-shot agents based on GPT Zhou et al. (2023); Long et al. (2023) have been proposed. However, they suffer from several limitations. For example, NavGPT Zhou et al. (2023) has limited performance and relies on a two-stage language-only system. DiscussNav Long et al. (2023) introduces a sequential multi-experts system to discuss and summarize various information and fuses five repeated predictions to improve performance. Some of their designs limit the agent’s capability to only address step-by-step instructions in the R2R dataset Anderson et al. (2018b), and have not been validated on other styles of instructions (e.g., REVERIE Qi et al. (2020b)). Besides, these agents are limited to local exploration as they can only reason and make decisions within adjacent navigable points. In this paper, we propose a map-guided prompting method with adaptive path planning for global exploration, achieving impressive performance on both R2R and REVERIE.
Large Language Models (LLMs) LLMs OpenAI (2023a, b, c); Chiang et al. (2023); Touvron et al. (2023a, b); Anil et al. (2023) have demonstrated remarkable capabilities in multiple domains. Recently, LLM-based agents Huang et al. (2023); Mu et al. (2023); Ahn et al. (2022); Pan et al. (2023); Brohan et al. (2023); Schumann et al. (2023); Hu et al. (2023) have also attracted significant interest of the community. For example, VoxPoser Huang et al. (2023) utilizes LLM and vision-language models to extract affordances and constraints, which enables motion planners to generate trajectories for manipulation. LangNav Pan et al. (2023) employs LLMs for navigation, but it merely utilizes GPT-4 OpenAI (2023a) to synthesize some data and performs fine-tuning using Llama2 Touvron et al. (2023b) as the backbone, rather than directly employing LLM as a zero-shot agent. In fact, the application of LLM-based zero-shot agents in navigation tasks is still limited, and how to prompt LLMs (including multimodal GPT-4V) to activate global thinking and planning abilities required by navigation task have yet to be explored Yang et al. (2023).
@@ 3.1 Single Expert Prompt System
As a comparison, our designed single expert prompt system has several features. (1) We eliminate the need for a separate design of an additional historical summary expert or instruction decomposing expert based on GPT models, which makes it convenient to incorporate both visual/textual inputs and additional information, such as maps. (2) Our navigation expert can utilize GPT-4V to make decisions directly based on visual observations in one stage. It can also take text descriptions as inputs and flexibly apply them in the two-stage GPT-4 system. (3) Simple yet efficient. In the R2R dataset, our two-stage system requires an average of 672 input tokens and 115 output tokens per step. In comparison, NavGPT utilizes three GPT experts, resulting in an average cost of 2,465 input tokens and 317 output tokens per step.
@@ 3.1.2 Fundamental Inputs
Our agent is equipped with an RGB camera to capture $M$ images of the environment. Unlike NavGPT Zhou et al. (2023) that gathers excessive environmental information, we prioritize the observations of navigable points, which can be formulated as $O_{t}=\{o_{t,i}\}_{i=1}^{N}$, where $N$ is the number of navigable viewpoints at step $t$. Each $o_{t,i}$ represents the observation towards a specific navigable point. These observations are original images for a one-stage system based on GPT-4V and can also be replaced with caption and detection results in a two-stage system, where we follow NavGPT Zhou et al. (2023) and utilize off-the-shelf vision models BLIP-2 Li et al. (2023) and Faster R-CNN Ren et al. (2015); Anderson et al. (2018a) to acquire scene description and object detection respectively. NavGPT also utilizes the bounding boxes provided by the REVERIE dataset to extract an additional object list of all the views, which is then used for the R2R experiment. We also utilize this object list and name it the “Surroundings”. However, we do not utilize this additional information in R2R. We only utilize it to enable the agent to determine whether to stop in the REVERIE dataset since we do not specify the directions of these surrounding objects.
@@ 4.2.1 Results on the Room-to-Room Dataset
Validation unseen set. We further compare the navigation performance between the proposed MapGPT and previous NavGPT on a larger validation unseen set with 11 scenes and 783 trajectories. As shown in Table 2, due to data distribution shift, the success rate of MapGPT is slightly lower (38.8% with GPT-4 and 43.7% with GPT-4V) compared to the results on the 72 scenes. Compared to NavGPT (GPT-4 based), MapGPT (GPT-4 based) exhibits a noticeable reduction in NE, leading by 4.8% in SR, and a substantial 15.6% improvement in OSR. Obviously, our proposed map-guided prompting with adaptive planning has raised the upper limit of agent navigation ability (OSR), thus increasing final success rate. However, due to information loss in the two-stage pipeline, MapGPT (GPT-4 based) tends to continue exploring when encountering insufficient description and has a weaker SPL performance. Nevertheless, the GPT-4V-based agent does not suffer from this issue, thus achieving 34.8% in SPL.
@@ 4.2.2 Results on the REVERIE Dataset
Benefiting from the flexible single-expert system and the utilization of a universal map-guided prompting and planning, we effortlessly apply MapGPT to the REVERIE dataset with different instruction styles. Considering the API costs and easier comparison for future work, we randomly sample 500 instructions from the validation unseen set for the zero-shot setting of the REVERIE benchmark.
@@
Backtracking ratio. We also compare the backtracking ratio of NavGPT and our MapGPT on REVERIE. MapGPT experiences at least one instance of backtracking in 49% of cases, and among them, there is an 80% probability of successfully correcting its erroneous navigation path at least once. As a comparison, NavGPT, which does not utilize any maps and planning, performs poorly on this metric. Only 18% of cases involve backtracking to visited nodes, and among these cases, 53% of instances successfully correct the previous exploration error at least once.
@@ 5 Conclusion
Despite the significant performance gap between MapGPT and models based on pre-training and fine-tuning, zero-shot VLN still holds significant research value. It requires no data training, thus demonstrating great potential in terms of generalization and sim-to-real transfer, as GPT’s pre-training corpus contains a large amount of real-world image data, while data from simulators is comparatively limited. However, MapGPT is only experimented within a simulator that incorporates certain ideal assumptions, which is a limitation. Developing GPT-based agents directly in the real world and addressing various real-world challenges would be a meaningful future direction.
@@ A.1 Prompts
For a one-stage system based on GPT-4V, we directly require the agent to refer to images corresponding to various places. At the end of all prompts, we input the image IDs and original images in an interleaved format, such as “Image 0: <Img0> Image 1: <Img1> …”. When used in the action space, an action $a_{t}$ could be “C. turn around to Place 2 which is corresponding to Image 2”. Prompts in history and maps also employ similar templates.
@@ A.2 Implementation Details
We have built the first GPT-4V-based OpenAI (2023b, c) agent in the VLN field, directly processing multimodal inputs in one stage. For a fair comparison with previous methods specifically designed for GPT-4, we have also implemented a two-stage system, where we follow NavGPT Zhou et al. (2023) and utilize BLIP-2 Li et al. (2023) to provide a caption for the observation, and employ Faster R-CNN Ren et al. (2015) to detect existing objects. Our core contributions, namely map-guided prompting and adaptive path planning, can be applied to both of these systems. To adapt our MapGPT to the REVERIE dataset, we only make some simple yet necessary modifications, which demonstrates that our MapGPT is more unified in the VLN task.

### v3

@@ Abstract
Embodied agents equipped with GPT as their brains have exhibited extraordinary decision-making and generalization abilities across various tasks. However, existing zero-shot agents for vision-and-language navigation (VLN) only prompt GPT-4 to select potential locations within localized environments, without constructing an effective “global-view” for the agent to understand the overall environment. In this work, we present a novel map-guided GPT-based agent, dubbed MapGPT, which introduces an online linguistic-formed map to encourage global exploration. Specifically, we build an online map and incorporate it into the prompts that include node information and topological relationships, to help GPT understand the spatial environment. Benefiting from this design, we further propose an adaptive planning mechanism to assist the agent in performing multi-step path planning based on a map, systematically exploring multiple candidate nodes or sub-goals step by step. Extensive experiments demonstrate that our MapGPT is applicable to both GPT-4 and GPT-4V, achieving state-of-the-art zero-shot performance on R2R and REVERIE simultaneously ($\sim$10and $\sim$12improvements in SR), and showcasing the newly emergent global thinking and path planning abilities of the GPT.
@@
Jiaqi Chen1        Bingqian Lin2        Ran Xu3         Zhenhua Chai3 Xiaodan Liang2       Kwan-Yee K. Wong1 1The University of Hong Kong 2Shenzhen Campus of Sun Yat-sen University 3Meituan Project: https://chen-judge.github.io/MapGPT/
@@ 1 Introduction
Large language models (LLMs) Touvron et al. (2023a, b); Chowdhery et al. (2022); Anil et al. (2023) have demonstrated strong performance in various domains. As the most powerful LLMs, the GPT series models Brown et al. (2020); OpenAI (2023a, b, c) can even serve as the brains of embodied agents Huang et al. (2023); Mu et al. (2023); Ahn et al. (2022), enabling them to engage in explicit thinking and decision-making process. Moreover, these GPT-based agents are typically zero-shot or few-shot, eliminating the burdensome tasks of data annotation and model training, and they also demonstrate remarkable sim-to-real abilities.
@@
However, such zero-shot VLN agents face some challenges. Firstly, these methods are developed for a multi-expert pure language system based on GPT-4, necessitating the conversion of visual observations into text and multiple rounds of text summarizations, which inevitably lead to information loss and hinder the application to multimodal LLMs (e.g., GPT-4V). More importantly, these agents make decisions based solely on the observations of the local environment. As shown in Figure 1 (left), given only the local action space, when an agent realizes that it has engaged in an erroneous exploration, it can only continue to explore surrounding environment aimlessly.
@@ 2 Related Work
Vision-and-Language Navigation (VLN) As a representative multi-modal embodied AI task, VLN requires an agent to combine human instructions and visual observations to navigate and locate targets in real-world scenes. Previous learning-based approaches Wang et al. (2019); Ma et al. (2019); Deng et al. (2020); Qi et al. (2020a) proposed various model architectures and trained their models on domain-specific datasets. Besides, pretrained models Hong et al. (2021); Chen et al. (2021b, 2022); Qiao et al. (2022); Guhur et al. (2021); An et al. (2022); Lin et al. (2022); Qiao et al. (2023); Wang et al. (2023); Pan et al. (2023) have been widely applied to produce better multi-modal representations. Recently, to address the reliance on domain-specific data and the possible sim-to-real gap, some zero-shot agents based on GPT Zhou et al. (2023); Long et al. (2023) have been proposed. However, they suffer from several limitations. For example, NavGPT Zhou et al. (2023) has limited performance and relies on a two-stage language-only system. DiscussNav Long et al. (2023) introduces a sequential multi-experts system to discuss and summarize various information and fuses five repeated predictions to improve performance. Some of their designs limit the agent’s capability to only address step-by-step instructions in the R2R dataset Anderson et al. (2018b), and have not been validated on other styles of instructions (e.g., REVERIE Qi et al. (2020b)). Besides, these agents are limited to local exploration as they can only reason and make decisions within adjacent navigable points. In this paper, we propose a map-guided prompting method with adaptive path planning for global exploration, achieving impressive performance on both R2R and REVERIE.
Large Language Models (LLMs) LLMs OpenAI (2023a, b, c); Chiang et al. (2023); Touvron et al. (2023a, b); Anil et al. (2023) have demonstrated remarkable capabilities in multiple domains. Recently, LLM-based agents Huang et al. (2023); Mu et al. (2023); Ahn et al. (2022); Pan et al. (2023); Brohan et al. (2023); Schumann et al. (2023); Hu et al. (2023); Lin et al. (2024) have also attracted significant interest of the community. For example, VoxPoser Huang et al. (2023) utilizes LLM and vision-language models to extract affordances and constraints, which enables motion planners to generate trajectories for manipulation. LangNav Pan et al. (2023) employs LLMs for navigation, but it merely utilizes GPT-4 OpenAI (2023a) to synthesize some data and performs fine-tuning using Llama2 Touvron et al. (2023b) as the backbone, rather than directly employing LLM as a zero-shot agent. In fact, the application of LLM-based zero-shot agents in navigation tasks is still limited, and how to prompt LLMs (including multimodal GPT-4V) to activate global thinking and planning abilities required by navigation task have yet to be explored Yang et al. (2023).
@@ 3.1 Single Expert Prompt System
Compared with previous works, our proposed single expert prompt system has several unique features. (1) We eliminate the need for a separate design of an additional historical summary expert or instruction decomposing expert based on GPT models, which makes it convenient to incorporate both visual/textual inputs and additional information, such as maps. (2) Our navigation expert can utilize GPT-4V to make decisions directly based on visual observations in one stage. It can also take text descriptions as inputs and flexibly apply them in the two-stage GPT-4 system. (3) Simple yet efficient. In the R2R dataset, our two-stage system requires an average of 672 input tokens and 115 output tokens per step. In comparison, NavGPT utilizes three GPT experts, resulting in an average cost of 2,465 input tokens and 317 output tokens per step.
@@ 3.1.2 Fundamental Inputs
Our agent is equipped with an RGB camera to capture $M$ images of the environment. Unlike NavGPT Zhou et al. (2023) that gathers excessive environmental information, we prioritize the observations of navigable points, which can be formulated as $O_{t}=\{o_{t,i}\}_{i=1}^{N}$, where $N$ is the number of navigable viewpoints at step $t$. Each $o_{t,i}$ represents the observation towards a specific navigable point. These observations are original images for a one-stage system based on GPT-4V and can also be replaced with caption and detection results in a two-stage system, where we follow NavGPT Zhou et al. (2023) and utilize off-the-shelf vision models BLIP-2 Li et al. (2023a) and Faster R-CNN Ren et al. (2015); Anderson et al. (2018a) to acquire scene description and object detection respectively. NavGPT also utilizes the bounding boxes provided by the REVERIE dataset to extract an additional object list of all the views, which is then used for the R2R experiment. We also utilize this object list and name it the “Surroundings”. However, we do not utilize this additional information in R2R. We only utilize it to enable the agent to determine whether to stop in the REVERIE dataset since we do not specify the directions of these surrounding objects.
@@ 4.2.1 Results on the Room-to-Room Dataset
Validation unseen set. We further compare the navigation performance between the proposed MapGPT and previous NavGPT on a larger validation unseen set with 11 scenes and 783 trajectories. As shown in Table 2, due to the distribution difference, the success rate of MapGPT is slightly lower (38.8% with GPT-4 and 43.7% with GPT-4V) compared to the results on the 72 scenes. Compared to NavGPT (GPT-4 based), MapGPT (GPT-4 based) exhibits a noticeable reduction in NE, leading by 4.8% in SR, and a substantial 15.6% improvement in OSR. Obviously, our proposed map-guided prompting with adaptive planning has raised the upper limit of agent navigation ability (OSR), thus increasing final success rate. However, due to information loss in the two-stage pipeline, MapGPT (GPT-4 based) tends to continue exploring when encountering insufficient description and has a weaker SPL performance. Nevertheless, the GPT-4V-based agent does not suffer from this issue, thus achieving 34.8% in SPL.
@@ 4.2.2 Results on the REVERIE Dataset
Benefiting from the flexible single-expert system and the utilization of a universal map-guided prompting and planning, we effortlessly apply MapGPT to the REVERIE dataset with different instruction styles. Considering the API costs and easier comparison for future work, we randomly sample 500 instructions from the validation unseen set for the zero-shot setting REVERIE benchmark.
@@
Backtracking ratio. We analyze the trajectory of each case, and if there are repeated visits to previously visited locations, we consider it as backtracking and calculate the ratio of such occurrences in both NavGPT and MapGPT for comparison on the REVERIE dataset. MapGPT experiences at least one instance of backtracking in 49% of cases, and among them, there is an 80% probability of successfully correcting its erroneous navigation path at least once. As a comparison, NavGPT, which does not utilize any maps and planning, performs poorly on this metric. Only 18% of cases involve backtracking to visited nodes, and among these cases, 53% of instances successfully correct the previous exploration error at least once.
@@ 5 Conclusion
Despite the significant performance gap between MapGPT and models based on pre-training and fine-tuning, zero-shot VLN still holds significant research value. GPT’s pre-training corpus contains a large amount of real-world image data, thus demonstrating great potential in terms of generalization and sim-to-real transfer. However, MapGPT is only experimented within a simulator that incorporates certain ideal assumptions. Developing LLM-based agents directly in the real world and addressing various real-world challenges would be a meaningful future direction.
Acknowledgements
This work was supported in part by CAAI-Huawei MindSpore Open Fund.
@@ A.1 Prompts
All of the environmental information collected at each step (including observations, history, maps, etc.) will be incorporated into the user messages of GPT.
@@
For a one-stage system based on GPT-4V, we directly require the agent to refer to images corresponding to various places. At the beginning of the user messages, we input the image IDs and observed images in an interleaved format, such as “Image 0: <Img0> Image 1: <Img1> …”. Therefore, these images can be directly referenced in subsequent prompts. For example, when used in the action space, an action $a_{t}$ could be “C. turn around to Place 2 which is corresponding to Image 2”. Prompts in history and maps also employ similar templates.
@@ A.2 Implementation Details
We conduct experiments in the Matterport3D simulator Chang et al. (2017), which provides a discrete navigation environment with predefined navigable viewpoints. At each viewpoint, the agent can obtain visual observations and some connected navigable candidate viewpoints which are incorporated into prompts for GPT. Once GPT has selected one of these candidates by predicting the corresponding label in the prompts, we can convert it into the candidate’s ID which can be executed in the simulator and teleport the agent to the selected viewpoint.
In this work, we have built the first GPT-4V-based OpenAI (2023b, c) agent in the VLN field, directly processing multimodal inputs in one stage. For a fair comparison with previous methods specifically designed for GPT-4 OpenAI (2023a), we have also implemented a two-stage system, where we follow NavGPT Zhou et al. (2023) and utilize BLIP-2 Li et al. (2023a) to provide a caption for the observation, and employ Faster R-CNN Ren et al. (2015) to detect existing objects. Our core contributions, namely map-guided prompting and adaptive path planning, can be applied to both of these systems. To adapt our MapGPT to the REVERIE dataset, we only make some simple yet necessary modifications, which demonstrates that our MapGPT is more unified in the VLN field.